# Chiral cat state generation via the Sagnac-Fizeau effect

### v1

@@ III.3 The Wigner functions of the exact states
Here the measurement probabilities $P_{\pm}(t)$ are given by Eq. (19), and $\mathcal{L}^{\epsilon=\mathrm{CW},\mathrm{CCW}}_{n,l}(\chi)$ is defined as $\mathcal{L}_{n,l}^{\epsilon}(\chi)={}_{\epsilon}\langle n|D(\chi)|l\rangle_{\epsilon}$, which is given by Oliveira1990
@@
In Fig. 5, we show the exact Wigner functions $\mathcal{W}^{(\pm)}_{\mathrm{CW,CCW}}(\chi)$ at time $t_{s}=69.11J^{-1}$. Figures 5(a) and 5(b) show clear interference fringes for the CW mode under the full Hamiltonian, which match well with the approximate solutions in Figs. 2(a) and 2(b). This indicates that the approximation is valid in describing the non-classical characteristics of the CW mode. For the CCW mode, however, the exact Wigner functions in Figs. 5(c) and 5(d) show small interference fringes that are not presented in the approximate solutions. These small fringes appear because the components $\alpha_{\mathrm{CCW}}^{(-)}(t)$ and $\alpha_{\mathrm{CCW}}^{(+)}(t)$ in phase space are not completely overlapped (the Wigner function does not form a perfectly circular Gaussian peak), resulting in some interference effects. Despite this, we can still clearly distinguish the CW and CCW modes under the full Hamiltonian, demonstrating the feasibility for generating chiral cat states with our scheme.
@@ IV.1 Quantum master equation and its exact solution
In real physical systems, the dissipations inevitably affect the dynamics of the system. For the present cavity-QED system, the average thermal excitation numbers for the cavity field and the atom are negligibly small, then we can assume that both the cavity fields and the atom are connected to individual vacuum baths. By introducing the dissipation terms for both the cavity modes and the atom, the dynamics of the system is described by the quantum master equation
@@
To solve the quantum master equation, we expand the density matrix of the system using the Fock-state bases as follows
@@
with $\rho_{w,m,k,v,n,j}=\langle w|_{\mathrm{CW}}\langle m|_{\mathrm{CCW}}\langle k|\rho|v\rangle n\rangle_{\mathrm{CW}}|j\rangle_{\mathrm{CCW}}$. To numerically solve these equations of motion for the matrix density-matrix elements, we choose the initial state of the system as $|+\rangle|\alpha\rangle_{\mathrm{CW}}|\alpha\rangle_{\mathrm{CCW}}$, then the corresponding initial values of these density matrix elements are given by $\rho_{e,m,k,e,n,j}(0)=\rho_{e,m,k,g,n,j}(0)=\rho_{g,m,k,e,n,j}(0)=\rho_{g,m,k,g,n,j}(0)=e^{-2|\alpha|^{2}}\alpha^{m+k}(\alpha^{*})^{n+j}/2\sqrt{m!n!k!j!}$. With the numerical solutions, we can track the evolution of the density matrix and investigate the physical properties of the generated states.
Figure 6: The fidelities (a) $f(t)$, (b) $f_{+}(t)$ , and (c) $f_{-}(t)$ as functions of the scaled evolution time $Jt$ for different cavity-field decay rates $\kappa/J=0.0025$, $0.005$, and $0.01$. The decay rate of the atom is $\gamma/J=0.005$. Other parameters used are the same as those in Fig. 2.
@@ IV.2 Fidelities between the approximate and exact states in the open-system case
where we introduce the variables $\Xi_{m,k,n,j}^{(\pm)}(t)=\rho_{e,m,k,e,n,j}(t)+\rho_{g,m,k,g,n,j}(t)\pm\rho_{e,m,k,g,n,j}(t)\pm\rho_{g,m,k,e,n,j}(t)$, and the measurement probabilities
@@
In Fig. 6, we show the fidelities $f(t)$ and $f_{\pm}(t)$ versus the scaled evolution time $Jt$ in the open-system case. Compared to the closed-system case in Fig. 4, the presence of system dissipation accelerates the decay of the fidelities and reduces oscillations. Specifically, increasing the cavity-field decay rate from $\kappa/J=0.0025$ to $0.005$ and $0.01$ leads to a more rapid fidelity reduction, as shown by the green solid, red dashed, and blue dotted curves, respectively. This trend is consistent for $f(t)$ and $f_{\pm}(t)$. Moreover, the results show that at time $t_{s}=69.11J^{-1}$, the fidelities $f(t_{s})$, $f_{+}(t_{s})$, and $f_{-}(t_{s})$ take the values 0.76, 0.77, and 0.76 for $\kappa/J=0.0025$; 0.66, 0.67, and 0.66 for $\kappa/J=0.005$; and 0.52, 0.52, and 0.52 for $\kappa/J=0.01$, respectively. These results show that larger decay rates lead to reduced fidelity, indicating the degression of quantum state generation by the dissipations in the open-system case.
Figure 7: The measurement probabilities (a) $p_{+}(t)$ and (b) $p_{-}(t)$ given in Eq. (27) versus the scaled evolution time $Jt$ for different cavity-field decay rates $\kappa/J=0.0025$, $0.005$, and $0.01$ in the open-system case. The parameters used are the same as those in Fig. 6.
@@ IV.3 The Wigner functions of the generated states in the open-system case
Figure 8: The Wigner functions $\mathcal{\tilde{W}}^{(\pm)}_{\mathrm{CW,CCW}}(\chi)$ corresponding to the states $\rho^{(\pm)}_{\epsilon=\mathrm{CW},\mathrm{CCW}}=\mathrm{Tr}_{\bar{\epsilon}}(\rho^{(\pm)})$, where $\rho^{(\pm)}$ are the states defined in Eq. (26). These Wigner functions are shown at the time $t_{s}=\pi/2\zeta_{\mathrm{CW}}$ for various cavity-field decay rates: (a)-(d) $\kappa/J=0.0025$, (e)-(h) $\kappa/J=0.005$, and (i)-(l) $\kappa/J=0.01$. Other parameters used are the same as those in Fig. 6.
@@
In Fig. 8, we show the Wigner functions $\mathcal{\tilde{W}}^{(\pm)}_{\mathrm{CW,CCW}}(\chi)$ of the generated states $\rho^{(\pm)}_{\epsilon=\mathrm{CW},\mathrm{CCW}}$ corresponding to different cavity-field decay rates. Compared to the closed-system results in Figs. 5(a) and 5(b), we observe that increasing the cavity-field decay rate reduces quantum coherence. For both the CW and CCW modes, when the decay rate takes $\kappa/J=0.0025$ [Figs. 8(a)-8(b)], $\kappa/J=0.005$ [Figs. 8(e)-8(h)], and $\kappa/J=0.01$ [Figs. 8(i)-8(l)], we find that the interference fringes in the corresponding Wigner functions become weaker for a larger $\kappa$. Despite the reduction of quantum coherence due to dissipation, the CW mode remains in a cat state, while the CCW mode remains in a coherent state. This confirms that the chiral-cat-state generation scheme is robust even in the open-system cases.
@@ V Discussions and conclusion
To analyze the experimental feasibility of our scheme, we present some discussions concerning the experimental setups and the relevant parameters. Provided the use of spinning resonator, we focus on the microring resonator system coupled to a two-level system Kavokin2007 . Currently, the strong coupling between a single atom and a monolithic microresonator has been observed DayanSci2008 ; Aokiat2006 , with an effective coupling strength of $J/2\pi=50\pm 12~{}\mathrm{MHz}$ Aokiat2006 . Moreover, the experimentally feasible parameters for the microresonator are $Q=6.0\times 10^{9}$, $R=1.1$ mm, $n_{r}=1.4$, $\lambda=1550$ nm, and $\Omega=6.6$ kHz MaayaniNa2018 . These parameters result in a cavity-field decay rate $\kappa\approx 0.2$ MHz. The ratio of the cavity-field decay rate to the coupling strength used is of the same order as the experimental parameters. In our numerical simulations, the Sagnac-Fizeau frequency shift used is $\Delta_{\mathrm{sag}}=11J$, which corresponds to an angular velocity of $\Omega\approx\Delta_{\mathrm{sag}}c/[n_{r}R\omega_{c}(1-n_{r}^{-2})]=\Delta_{\mathrm{sag}}\lambda/[2\pi n_{r}R(1-n_{r}^{-2})]\approx 2\pi\times 136.9$ kHz for the above experimental parameters. We point out that this rotation angular velocity can be decreased by choosing a smaller coupling strength $J$. Recently, an angular velocity of $6.6$ kHz has been reported. These analyses indicate that the suggested parameters should be accessible with the near future experimental technique. We also note that, even at lower angular velocities, we can still generate chiral cat states, where one mode remains in a coherent state, while the other mode is in a less distinguishable cat state.
In conclusion, we have proposed a reliable method for creating chiral cat states in a cavity-QED system consisting of a spinning resonator coupled to a two-level atom in the dispersive-coupling regime. We have shown that a cat state in the CW mode and a coherent state in the CCW mode can be successfully generated. By calculating the Wigner function, we have characterized quantum coherence properties of these generated cat states. In addition, we have examined the influence of the system dissipations on chiral cat state generation. Our work will contribute to the development of chiral quantum optics and the explorement of chiral optical devices.
@@ Acknowledgements.
J.-Q.L. was supported in part by National Natural Science Foundation of China (Grants No. 12175061, No. 12247105, No. 11935006, and No. 12421005), National Key Research and Development Program of China (Grant No. 2024YFE0102400), and Hunan Provincial Major Sci-Tech Program (Grant No. 2023ZJ1010). H.J. is supported by the Natural Science Foundation of China (NSFC) (Grant No. 11935006), the Science and Technology Innovation Program of Hunan Province (Grant No. 2020RC4047), National Key R&D Program of China (Grant No. 2024YFE0102400), and Hunan provincial major Sci-Tech program (Grant No. 2023ZJ1010). L.-M.K. is supported by the Natural Science Foundation of China (NSFC) (Grant Nos. 12247105, 12175060, 11935006, 12421005), Hunan Provincial Major Sci-Tech Program (Grant No. 2023ZJ1010), and XJ-Lab key project (Grant No. 23XJ02001). Y.-H.L. was supported in part by Hunan Provincial Postgraduate Research and Innovation project (Grant No. CX20240530).

### v2

@@ III.3 Wigner functions of the exact states
Here the measurement probabilities $P_{\pm}(t)$ are given by Eq. (19), and $\mathcal{L}^{\epsilon=\mathrm{CW},\mathrm{CCW}}_{n,l}(\chi)$ is defined as $\mathcal{L}_{n,l}^{\epsilon}(\chi)={}_{\epsilon}\langle n|D(\chi)|l\rangle_{\epsilon}$, which is given by [88]
@@
In Fig. 5, we show the exact Wigner functions $\mathcal{W}^{(\pm)}_{\mathrm{CW,CCW}}(\chi)$ at time $t_{s}=69.11J^{-1}$. Figures 5(a) and 5(b) show clear interference fringes for the CW mode under the full Hamiltonian, which match well with the approximate solutions in Figs. 2(a) and 2(b). This indicates that the approximation is valid in describing the nonclassical characteristics of the CW mode. For the CCW mode, however, the exact Wigner functions in Figs. 5(c) and 5(d) show small interference fringes that are not presented in the approximate solutions. These small fringes appear because the components $\alpha_{\mathrm{CCW}}^{(-)}(t)$ and $\alpha_{\mathrm{CCW}}^{(+)}(t)$ in phase space are not completely overlapped (the Wigner function does not form a perfectly circular Gaussian peak), resulting in some interference effects. Despite this, we can still clearly distinguish the CW and CCW modes under the full Hamiltonian, demonstrating the feasibility for generating chiral cat states with our scheme.
@@ IV.1 Quantum master equation and its exact solution
In real physical systems, the dissipations inevitably affect the dynamics of the system. For the present cavity-QED system, the average thermal excitation numbers for the cavity field and the atom are negligibly small; then we can assume that both the cavity fields and the atom are connected to individual vacuum baths. By introducing the dissipation terms for both the cavity modes and the atom, the dynamics of the system is described by the quantum master equation
@@
To solve the quantum master equation, we expand the density matrix of the system using the Fock-state bases as follows:
@@
with $\rho_{w,m,k,v,n,j}=\langle w|_{\mathrm{CW}}\langle m|_{\mathrm{CCW}}\langle k|\rho|v\rangle n\rangle_{\mathrm{CW}}|j\rangle_{\mathrm{CCW}}$. To numerically solve these equations of motion for the matrix density-matrix elements, we choose the initial state of the system as $|+\rangle|\alpha\rangle_{\mathrm{CW}}|\alpha\rangle_{\mathrm{CCW}}$; then the corresponding initial values of these density matrix elements are given by $\rho_{e,m,k,e,n,j}(0)=\rho_{e,m,k,g,n,j}(0)=\rho_{g,m,k,e,n,j}(0)=\rho_{g,m,k,g,n,j}(0)=e^{-2|\alpha|^{2}}\alpha^{m+k}(\alpha^{*})^{n+j}/2\sqrt{m!n!k!j!}$. With the numerical solutions, we can track the evolution of the density matrix and investigate the physical properties of the generated states.
@@ IV.2 Fidelities between the approximate and exact states in the open-system case
Figure 6: Fidelities (a) $f(t)$, (b) $f_{+}(t)$ , and (c) $f_{-}(t)$ as functions of the scaled evolution time $Jt$ for different cavity-field decay rates $\kappa/J=0.0025$, $0.005$, and $0.01$. The decay rate of the atom is $\gamma/J=0.005$. Other parameters used are the same as those in Fig. 2.
@@
where we introduce the variables $\Xi_{m,k,n,j}^{(\pm)}(t)=\rho_{e,m,k,e,n,j}(t)+\rho_{g,m,k,g,n,j}(t)\pm\rho_{e,m,k,g,n,j}(t)\pm\rho_{g,m,k,e,n,j}(t)$ and the measurement probabilities
@@
Figure 7: Measurement probabilities (a) $p_{+}(t)$ and (b) $p_{-}(t)$ given in Eq. (27) versus the scaled evolution time $Jt$ for different cavity-field decay rates $\kappa/J=0.0025$, $0.005$, and $0.01$ in the open-system case. The parameters used are the same as those in Fig. 6.
In Fig. 6, we show the fidelities $f(t)$ and $f_{\pm}(t)$ versus the scaled evolution time $Jt$ in the open-system case. Compared to the closed-system case in Fig. 4, the presence of system dissipation accelerates the decay of the fidelities and reduces oscillations. Specifically, increasing the cavity-field decay rate from $\kappa/J=0.0025$ to $0.005$ and $0.01$ leads to a more rapid fidelity reduction, as shown by the green solid, red dashed, and blue dotted curves, respectively. This trend is consistent for $f(t)$ and $f_{\pm}(t)$. Moreover, the results show that, at time $t_{s}=69.11J^{-1}$, the fidelities $f(t_{s})$, $f_{+}(t_{s})$, and $f_{-}(t_{s})$ take the values 0.76, 0.77, and 0.76 for $\kappa/J=0.0025$, 0.66, 0.67, and 0.66 for $\kappa/J=0.005$, and 0.52, 0.52, and 0.52 for $\kappa/J=0.01$, respectively. These results show that larger decay rates lead to reduced fidelity, indicating the degression of quantum state generation by the dissipations in the open-system case.
@@ IV.3 The Wigner functions of the generated states in the open-system case
Figure 8: Wigner functions $\mathcal{\tilde{W}}^{(\pm)}_{\mathrm{CW,CCW}}(\chi)$ corresponding to the states $\rho^{(\pm)}_{\epsilon=\mathrm{CW},\mathrm{CCW}}=\mathrm{Tr}_{\bar{\epsilon}}(\rho^{(\pm)})$, where $\rho^{(\pm)}$ are the states defined in Eq. (26). These Wigner functions are shown at the time $t_{s}=\pi/(2\zeta_{\mathrm{CW}})$ for various cavity-field decay rates: (a)--(d) $\kappa/J=0.0025$, (e)--(h) $\kappa/J=0.005$, and (i)--(l) $\kappa/J=0.01$. Other parameters used are the same as those in Fig. 6.
In Fig. 8, we show the Wigner functions $\mathcal{\tilde{W}}^{(\pm)}_{\mathrm{CW,CCW}}(\chi)$ of the generated states $\rho^{(\pm)}_{\epsilon=\mathrm{CW},\mathrm{CCW}}$ corresponding to different cavity-field decay rates. Compared to the closed-system results in Figs. 5(a) and 5(b), we observe that increasing the cavity-field decay rate reduces quantum coherence. For both the CW and CCW modes, when the decay rate takes $\kappa/J=0.0025$ [Figs. 8(a)--8(d)], $\kappa/J=0.005$ [Figs. 8(e)--8(h)], and $\kappa/J=0.01$ [Figs. 8(i)--8(l)], we find that the interference fringes in the corresponding Wigner functions become weaker for a larger $\kappa$. Despite the reduction of quantum coherence due to dissipation, the CW mode remains in a cat state, while the CCW mode remains in a coherent state. This confirms that the chiral-cat-state generation scheme is robust even in the open-system cases.
@@ V   Discussions and conclusion
To analyze the experimental feasibility of our scheme, we present some discussions concerning the experimental setups and the relevant parameters. Provided the use of spinning resonator, we focus on the microring resonator system coupled to a two-level system [89]. Currently, the strong coupling between a single atom and a monolithic microresonator has been observed [91, 90], with an effective coupling strength of $J/2\pi=50\pm 12~{}\mathrm{MHz}$ [90]. Moreover, the experimentally feasible parameters for the microresonator are $Q=6.0\times 10^{9}$, $R=1.1$ mm, $n_{r}=1.4$, $\lambda=1550$ nm, and $\Omega=6.6$ kHz [72]. These parameters result in a cavity-field decay rate $\kappa\approx 0.2$ MHz. The ratio of the cavity-field decay rate to the coupling strength used is of the same order as the experimental parameters. In our numerical simulations, the Sagnac-Fizeau frequency shift used is $\Delta_{\mathrm{sag}}=11J$, which corresponds to an angular velocity of $\Omega\approx\Delta_{\mathrm{sag}}c/[n_{r}R\omega_{c}(1-n_{r}^{-2})]=\Delta_{\mathrm{sag}}\lambda/[2\pi n_{r}R(1-n_{r}^{-2})]\approx 2\pi\times 136.9$ kHz for the above experimental parameters. We point out that this rotation angular velocity can be decreased by choosing a smaller coupling strength $J$. Recently, an angular velocity of $6.6$ kHz has been reported. These analyses indicate that the suggested parameters should be accessible with the near future experimental technique. We also note that, even at lower angular velocities, we can still generate chiral cat states, where one mode remains in a coherent state, while the other mode is in a less distinguishable cat state.
In conclusion, we have proposed a reliable method for creating chiral cat states in a cavity-QED system consisting of a spinning resonator coupled to a two-level atom in the dispersive-coupling regime. We have shown that a cat state in the CW mode and a coherent state in the CCW mode can be successfully generated. By calculating the Wigner function, we have characterized quantum coherence properties of these generated cat states. In addition, we have examined the influence of the system dissipations on chiral-cat-state generation. Our work will contribute to the development of chiral quantum optics and the exploration of chiral optical devices.
@@ Acknowledgements.
J.-Q.L. was supported in part by the National Natural Science Foundation of China (Grants No. 12175061, No. 12247105, No. 11935006, and No. 12421005), the National Key Research and Development Program of China (Grant No. 2024YFE0102400), and the Hunan Provincial Major Sci-Tech Program (Grant No. 2023ZJ1010). H.J. is supported by the Natural Science Foundation of China (NSFC) (Grant No. 11935006), the Science and Technology Innovation Program of Hunan Province (Grant No. 2020RC4047), the National Key R&D Program of China (Grant No. 2024YFE0102400), and the Hunan provincial major Sci-Tech program (Grant No. 2023ZJ1010). L.-M.K. is supported by the Natural Science Foundation of China (NSFC) (Grants No. 12247105, No. 12175060, No. 11935006, No. 12421005), the Hunan Provincial Major Sci-Tech Program (Grant No. 2023ZJ1010), and the XJ-Lab key project (Grant No. 23XJ02001). Y.-H.L. was supported in part by the Hunan Provincial Postgraduate Research and Innovation project (Grant No. CX20240530).